10. Calculate ∆Hrxn for the reaction: CH4(g) + 4Cl2 → CCl4(g) + 4HCl(g) Given that: C(s) + 2H2(g) → CH4(g) C(s) + 2Cl2(g) → CCl4(g) H2(g) + Cl2(g) → 2HCl(g) ∆H = -74.6kJ ∆H = -95.7kJ ∆H = -92.3kJ
Answer #1

e(s)+ 2H2(g) + AH = -74.6 KJ .... C(s) + 241269) → eeld (9) - 4H= -95.745.- H2(g) + Cl249)-> 2H2Cg), AH = -92.3KJ.-3 2x eqn 3